McDougal Company issued a $500,000, 5-year, 9%APR bond dated January 1, 2018, on March 31, 2018 at par plus accrued interest. The bonds pay interest semiannually on June 30 and December 31. McDougal also incurred bond issue costs of $10,000.

Determine the Cash Proceeds from the issue of this bond:

A. $511,250

B. $490,000

C. $500,000

D. $501,250
Answer #1

Cash proceeds from the issue of bond = 500000 + 500000x9%*3/12 = $511250

Net cash proceeds = 511250 - 10000 = $501250